### v3

# Robust functional PCA for relative data

Abstract
This paper introduces a robust approach to functional principal component analysis (FPCA) for relative data, particularly density functions. While recent papers have studied density data within the Bayes space framework, there has been limited focus on developing robust methods to effectively handle anomalous observations and large noise. To address this, we extend the Mahalanobis distance concept to Bayes spaces, proposing its regularized version that accounts for the constraints inherent in density data. Based on this extension, we introduce a new method, robust density principal component analysis (RDPCA), for more accurate estimation of functional principal components in the presence of outliers. The method’s performance is validated through simulations and real-world applications, showing its ability to improve covariance estimation and principal component analysis compared to traditional methods.
Keywords Relative data $\cdot$ Bayes spaces $\cdot$ Robust principal component analysis $\cdot$ Functional data analysis.
1 Introduction
Functional data analysis (FDA) has developed rapidly over the past decade (e.g., Ramsay and Silverman (2005)), yet many applied settings involve relative functional data—functions where the key information lies not in their absolute magnitude, but the dynamics or distributional pattern of the measured quantity carries relevant information. Examples include age-specific fertility or mortality rates, spectral curves in materials science, and normalized sensor signals, where data are typically rescaled and interpreted in terms of relative variation across a domain.
These data often share structural features with probability density functions (PDFs)—non-negativity, unit integral, and a relative (compositional) structure. While PDFs are a motivating case, the broader goal is to analyze functions where the shape or internal distribution carries the signal of interest, and scale is irrelevant or removed. For the simplicity of the notation, we refer to such functional data as densities, emphasizing that the probability density functions are just one example of the functional data we consider in this context.
To analyze such data, we use compositional data analysis (CoDA) extended to infinite dimensions, modeling functions as infinite-dimensional compositions in a Bayes space, a Hilbert space respecting their relative nature (Egozcue et al., 2006; van den Boogaart et al., 2010). Within this framework, simplicial functional PCA (SFPCA) (Hron et al., 2016) offers a dimension reduction tool tailored to relative functional data. However, SFPCA depends on the sample covariance operator and is sensitive to outliers, limiting its robustness in practice. To address this, robust alternatives using trimmed covariance estimators and center-outward orderings have been proposed. Depth measures such as spatial depth (Menafoglio et al., 2021) and others (Fraiman and Muniz, 2001; Febrero et al., 2008; López-Pintado and Romo, 2009; Arribas-Gil and Romo, 2014; Chakraborty and Chaudhuri, 2014) offer one route, while Mahalanobis-type distances—widely used in multivariate and functional settings (Secchi et al., 2013; Galeano et al., 2015; Ghiglietti et al., 2017; Berrendero et al., 2020)—provide another, such as the minimum regularized covariance trace (MRCT) estimator (Oguamalam et al., 2024). In this paper, we focus on the latter group of methods, extending the concept of regularized Mahalanobis distance for univariate functional data (later called RDMD) to the Bayes space. We then use this notion to construct a robust covariance estimator and a corresponding robust PCA method tailored for densities.
The rest of the paper is organized as follows: Section 2 discusses the concept of Bayes spaces and SFPCA and showcases potential problems that the outliers in the sample can cause. In Section 3 we propose a way of standardizing (whitening) relative functional data, thus extending and generalizing the notions of truncated- and $\alpha$-Mahalanobis distance from Galeano et al. (2015) and Berrendero et al. (2020), respectively, to Bayes Hilbert spaces. In Section 4, the novel, robust density PCA (RDPCA) method is introduced. Its performance is evaluated in a simulation study and real-data examples in Sections 5 and 6, respectively. Proofs of technical results are given in Supplement Section LABEL:supp:proofs.
2 Preliminaries
In traditional (multivariate) compositional data analysis, the emphasis is on the information found in the relative relationships between variables. This approach led to the development of CoDA based on the Aitchison geometry (Pawlowsky-Glahn et al., 2015). CoDA focuses on positive multivariate observations that are scale invariant, i.e. that can be arbitrarily rescaled equivalence classes of proportional vectors thus form the sample space of compositional data. Along with perturbation and powering operations these concepts form a linear space that operates on the sample space of compositions.
2.1 Bayes space
If we expand the horizon to infinite-dimensional compositions, we consider positive functions that are typically represented with unit integral constraint. Intuitively, such compositions can be thought of as density functions. Therefore, statistical concepts designed for unconstrained functional data need to be adapted, as they do not account for this feature, leading to potentially meaningless results when applied to densities (Delicado, 2011; Hron et al., 2016; Menafoglio et al., 2018). Consequently, to work with functional compositions, it is necessary to establish the fundamental concepts of the underlying space structure.
The foundational results were initially presented by Egozcue et al. (2006) and were subsequently extended by van den Boogaart et al. (2010), van den Boogaart et al. (2014), and Talská et al. (2020). The framework for probability density functions is based on a measurable space $(\Omega,\mathcal{A})$, typically set to $(\mathbb{R},\mathfrak{B}(\mathbb{R}))$, with $\mathfrak{B}(\mathbb{R})$ as the Borel $\sigma$-algebra on $\mathbb{R}$. Within this space, we consider the set of measures $\mu$ that are absolutely continuous with respect to a reference measure $\lambda$, here assumed to be the Lebesgue measure with compact support $I=[a,b]\subset\mathbb{R}$ and with length $\eta:=b-a$. While the authors of van den Boogaart et al. (2014) developed the theory of Bayes space for more general measures $\lambda$, in practice, commonly the Lebesgue measure is chosen. Furthermore, usually the measures $\mu$ are identified with their respective Radon-Nikodym derivative, density $f$. For these densities, a notion of equivalence in the Bayes sense for multiples is introduced, i.e $f=_{B}g$ if $f=c\cdot g$ for some $c\in\mathbb{R}.$ W.l.o.g. within these $(=_{B})$-equivalence classes, we take the element satisfying the unit integral constraint as the representative.
To equip the sample space of these density functions with a structure of a Hilbert space on a given bounded domain, Egozcue et al. (2006) followed the ideas grounding the Aitchison geometry, and extended its operations $(\oplus,\odot)$ (a.k.a. perturbation and powering), and the inner product $\langle\cdot,\cdot\rangle$ to the infinite-dimensional setting by defining
$(f\oplus g)(t)=_{B}f(t)g(t),\quad(\alpha\odot f)(t)=_{B}(f(t))^{\alpha},\quad\langle f,g\rangle_{B^{2}}=\frac{1}{2\eta}\int_{I}\int_{I}\text{log}\frac{f(t)}{f(s)}\text{log}\frac{g(t)}{g(s)}\mathrm{d}t\,\mathrm{d}s,$
for densities $f,g$ and $\alpha\in\mathbb{R}$. In this context, negative perturbation $(\ominus)$ is defined as $f\ominus g=f\oplus(-1\odot g)=_{B}f/g$. Endowing the subspace $\mathcal{B}^{2}(I)$ of positive densities on $I$ with a square-integrable logarithm (van den Boogaart et al., 2014) with this structure yields a separable Hilbert space. In practice, the centered logratio (clr) transformation
$\text{clr}(f)=\text{log}(f)-\frac{1}{\eta}\int_{I}(\text{log}(f))(t)\mathrm{d}t$ (1)
is used to directly work in a space equipped with a Euclidean structure. Observe that all clr transformed densities satisfy the zero integral constraint.
Denoting $L_{0}^{2}(I)\subset L^{2}(I)$ the set of all square integrable functions on $I$ that additionally integrate to zero, it can be shown that the clr transformation is an isomorphism between $\mathcal{B}^{2}(I)$ and $L_{0}^{2}(I)$, with the corresponding inverse transformation $\mathrm{clr}^{-1}(f)=_{B}\exp{(f)}$, for $f\in L_{0}^{2}(I)$. The clr transformation (1) enables thus the computation of perturbation, powering, and the inner product of densities through their transformed counterparts in the $L^{2}$ space. Let $f,g\in\mathcal{B}^{2}(I)$ and $\alpha,\beta\in\mathbb{R}$, then
$\langle f,g\rangle_{B^{2}}=~\langle\text{clr}(f),\text{clr}(g)\rangle_{L^{2}}~~\text{and}~~\text{clr}((\alpha\odot f)\oplus(\beta\odot g))=~\alpha\cdot\text{clr}(f)+\beta\cdot\text{clr}(g),$
making clr transformation an isometry.
2.2 Moments of random densities
A Bayes space is a Hilbert space. Therefore, we use the notion of the mean and the covariance in a general Hilbert space structure to specify and inspect the properties of the Bayes mean density $\mu_{B^{2},X}$ and Bayes covariance $C_{B^{2},X}$ of $X\in\mathcal{B}^{2}(I)$; see e.g. Kokoszka and Reimherr (2017) for more detail. For a random density $X\in\mathcal{B}^{2}(I)$, a Bayes mean density $\mu_{B^{2},X}$ is an object in $\mathcal{B}^{2}(I)$ satisfying
$\langle\mu_{B^{2},X},f\rangle_{B^{2}}=\mathbb{E}(\langle X,f\rangle_{B^{2}}),$ (2)
for every $f\in\mathcal{B}^{2}(I)$. Alternatively, the Frechet mean approach used in Petersen et al. (2022) can be used to obtain the same definition. Using the equivalent identity of covariance operators, we define the Bayes covariance operator $C_{B^{2},X}:\mathcal{B}^{2}(I)\to\mathcal{B}^{2}(I)$ of a random density $X\in\mathcal{B}^{2}(I)$ as a $\mathcal{B}^{2}(I)$ operator satisfying
$\langle C_{B^{2},X}f,g\rangle_{B^{2}}=\mathbb{E}\left(\langle X\ominus\mu_{B^{2},X},f\rangle_{B^{2}}\langle X\ominus\mu_{B^{2},X},g\rangle_{B^{2}}\right),$
for every $f,g\in\mathcal{B}^{2}(I)$. We write $C_{B^{2},X}=\mathbb{E}((X\ominus\mu_{B^{2},X})\otimes_{B^{2}}(X\ominus\mu_{B^{2},X})),$ where for $f,g\in\mathcal{B}^{2}(I)$ the outer product $f\otimes_{B^{2}}g:\mathcal{B}^{2}(I)\to\mathcal{B}^{2}(I)$ is the linear operator satisfying $(f\otimes_{B^{2}}g)(h)=\langle g,h\rangle_{B^{2}}f$, for $h\in\mathcal{B}^{2}(I)$. We require the density variable $X$ to possess finite fourth moment, i.e., $\mathbb{E}[\|X\|^{4}_{B^{2}}]<\infty$. Under this assumption, the covariance operator is a symmetric positive-definite Hilbert-Schmidt operator and admits the eigen-decomposition
$C_{B^{2},X}f=\sum_{j=1}^{+\infty}\lambda_{j}\langle f,\xi_{j}\rangle\xi_{j},$
where $(\lambda_{j},\xi_{j})$ are the eigenpairs of $C_{B^{2},X}$, conventionally ordered as $\lambda_{1}\geq\lambda_{2}\geq...\geq 0$. Note that all the eigenvalues are non-negative and, under the conditions above, $C_{B^{2},X}$ is trace-class, i.e., $\sum_{j=1}^{+\infty}\lambda_{j}<\infty$ (see, e.g., Horváth and Kokoszka (2012)). Lemma 2.1 gives the connection between the Bayes mean density $\mu_{B^{2},X}$ and covariance $C_{B^{2},X}$ and the “traditional" $L^{2}$-mean and the covariance of $\mathrm{clr}(X)$. The proof of Lemma 2.1 can be found in Section LABEL:supp:proofs of the supplement.
Lemma 2.1
Let $X\in\mathcal{B}^{2}(I)$ and $\mathrm{clr}(X)\in L^{2}(I)$ be its clr transformation. Then the following identities hold:
  • i)
    $\mathrm{clr}(\mu_{B^{2},X})=_{a.e}\mathbb{E}(\mathrm{clr}(X))$.
  • ii)
    $\mathrm{clr}(C_{B^{2},X}f)=_{a.e.}C_{\mathrm{clr}(X)}\mathrm{clr}(f)$, for every $f\in\mathcal{B}^{2}(I)$.
  • iii)
    Let $(\lambda_{i},\zeta_{i})$, $i\geq 1$ be the eigenpairs of $C_{B^{2},X}$. Then, $(\lambda_{i},\mathrm{clr}(\zeta_{i})/\int_{I}\mathrm{clr}(\zeta_{i}(t))\mathrm{d}t)$, $i\geq 1$ are the eigenpairs of $C_{\mathrm{clr}(X)}$.
2.3 Simplicial functional PCA
As in the multivariate principal component analysis, functional PCA aims to search for the main modes of variability within a dataset, see e.g. Ramsay and Silverman (2005). More recently, this concept was extended to the Bayes spaces setting in Hron et al. (2016). The authors coined the term simplicial functional PCA, referring to the simplex, which is the multivariate counterpart of the Bayes space. More precisely, given a data set $X_{1},\ldots,X_{n}\in\mathcal{B}^{2}(I)$, the problem of finding the first principal function $\zeta_{1}$ is defined as
[…]
where $\hat{\mu}_{B^{2},X}=\frac{1}{n}\odot\bigoplus_{i=1}^{n}X_{i}$ is the sample Bayes-mean. Subsequent components $\{\zeta_{i}\}_{i\geq 2}$ are then found by the same objective function under the additional orthogonality constraint $\langle\zeta_{i},\zeta_{j}\rangle_{B^{2}}=0$ for $j<i$. The principal functions can be found solving the eigenvalue problem
$\hat{C}_{B^{2},X}\left(\zeta_{j}\right)=\lambda_{j}\odot\zeta_{j},~j\geq 1$ (4)
where $\hat{C}_{B^{2},X}(Y)=\frac{1}{n}\odot\bigoplus_{i=1}^{n}\langle X_{i}\ominus\hat{\mu}_{B^{2},X},Y\rangle_{B^{2}}\odot(X_{i}\ominus\hat{\mu}_{B^{2},X})$ is the sample covariance operator in $\mathcal{B}^{2}(I)$ and $\lambda_{j},j\geq 1$, are its eigenvalues. As discussed in Hron et al. (2016), this problem can be solved in $L^{2}(I)$ by using the clr transformation; see also Lemma 2.1.
However, as previously emphasized, the eigendecomposition (4) and consequently also the optimization problem (3), are sensitive to the presence of outlying curves in the sample. Therefore, in order to achieve a robust estimation of the functional PCs, the underlying covariance will be based on a sub-sample consisting of the most central data points. The centrality of each observation will here be quantified by an adaptation of a regularized functional Mahalanobis distance for densities. For more insight into the notion of centrality/outlyingness for densities, see Remark 2.1.
Remark 2.1
In contrast to traditional FDA settings, identifying atypical curves in Bayes spaces requires rethinking outlier concepts. Because relative functions are compositional, local deviations affect the entire curve, and scaling, a key factor in conventional magnitude outliers, has no effect. Thus, outliers in this context reflect structural (shape) anomalies, such as unexpected redistribution of mass or signal across the domain. We detect such curves using the proposed concept of robust, regularized MD, flagging outliers based on the chosen quantile of its limiting distribution. We emphasize, however, that the goal of the present work is not to develop a method for outlier detection or to provide a taxonomy of outliers in relative functional data. Rather, we propose a novel approach to constructing robust PCA methods for relative functional data.
3 Notion of Mahalanobis distance for densities
In recent years, various attempts to extend the concept of multivariate Mahalanobis distance to $L^{2}$ spaces have been proposed. The primary challenge in this functional extension is the non-invertibility of the covariance operator in infinite-dimensional cases. For instance, for a fixed $p$, Galeano et al. (2015) defines the $p$-truncated Mahalanobis semidistance of $Y\in L^{2}(I)$ as a Mahalanobis distance of the corresponding optimal rank-$p$ projection; observe that it is indeed only a semidistance, as it lacks identifiability (zero truncated Mahalanobis distance does not necessarily imply a.s. zero process). Additionally, as argued in Berrendero et al. (2020), the partial sums corresponding to the Mahalanobis distance of optimal rank-$p$ projection diverge with $p$ and thus can depend heavily on the choice of $p$. Ghiglietti et al. (2017) proposed a modification of the approach presented in Galeano et al. (2015) to deal with the convergence issues. Berrendero et al. (2020) introduced the concept of $\alpha$-Mahalanobis distance, motivated by the Reproducing Kernel Hilbert Space (RKHS) associated with the underlying stochastic process that generates the data.
In a traditional multivariate setting, the Mahalanobis distance $M(\textbf{x},\textbf{y};\boldsymbol{\Sigma})$ between two random vectors $\textbf{x},\textbf{y}\in\mathbb{R}^{p}$ with respect to a regular covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{p\times p}$ can be thought of as a Euclidean distance between the two vectors, previously standardized (whitened) by the common covariance $M^{2}(\textbf{x},\textbf{y};\boldsymbol{\Sigma})=(\textbf{x}-\textbf{y})^{\prime}\boldsymbol{\Sigma}^{-1}(\textbf{x}-\textbf{y})=\|\boldsymbol{\Sigma}^{-1/2}\textbf{x}-\boldsymbol{\Sigma}^{-1/2}\textbf{y}\|^{2}.$ In that manner, the corresponding Mahalanobis norm of a random vector, i.e. the Mahalanobis distance of that random vector from its mean with respect to its covariance, equals the Euclidean norm of that random vector when standardized by its mean and the covariance. That is, $M(\textbf{x};\boldsymbol{\mu},\boldsymbol{\Sigma}):=M(\textbf{x},\boldsymbol{\mu};\boldsymbol{\Sigma})=\|\textbf{x}_{\mathrm{st}}\|$, where $\textbf{x}_{\mathrm{st}}:=\boldsymbol{\Sigma}^{-1/2}(\textbf{x}-\boldsymbol{\mu})$ is a solution to the linear problem $\boldsymbol{\Sigma}^{1/2}\textbf{x}_{\mathrm{st}}=\textbf{x}-\boldsymbol{\mu}$; for more details on whitening of multivariate data see e.g. (Friedman, 1987; Oja et al., 2016; Kessy et al., 2018).
In infinite-dimensional spaces like $L^{2}(I)$, the covariance is not invertible, implying that this latter linear problem is ill-posed. One of the recent proposals on how to address this problem was given in Oguamalam et al. (2024); Berrendero et al. (2020), using Tikhonov regularization. More specifically, given the regularization parameter $\alpha>0$, Oguamalam et al. (2024) define the $\alpha$-standardization of a random function $Y\in L^{2}(I)$ with mean $\mu$} and covariance operator $C$ as a solution to the optimization problem
$Y_{\mathrm{st}}^{\alpha}:=\operatorname*{arg\,min}_{Z\in L^{2}(I)}\left\{\|C^{1/2}Z-(Y-{\mu})\|^{2}+\alpha\|Z\|^{2}\right\}.$ (5)
Oguamalam et al. (2024) show that the $\alpha$-Mahalanobis distance is the Euclidean norm of the solution to (5). Regularization in (5) is also known as $L^{2}$-regularization due to penalizing the $L^{2}$-norm of the solution and is perhaps the most widely used penalization form. However, it is often advantageous to consider regularization using a more general regularization operator $L:L^{2}(I)\to L^{2}(I)$; see e.g. Morigi et al. (2007); Wang (2012).
Here, we extend the optimization problem in (5) to Bayes spaces, allowing for a broader class of regularization operators under general conditions.
3.1 Regularized standardization of density data
Standardization (whitening) is generally defined with respect to the mean and the covariance of the random object being standardized. For the sake of readability, we first discuss the regularized standardization, assuming the mean and the covariance are known.
Let, therefore, $X\in\mathcal{B}^{2}(I)$, with mean $\mu_{B^{2}}$ and covariance $C_{B^{2},X}$ as defined in Section 2. Without loss of generality, assume that $X$ has a constant mean (the neutral element for addition in $\mathcal{B}^{2}(I)$ is the density of the uniform distribution over $I$, i.e., a constant function).
Given the closed, densely defined operator $L:\mathcal{B}^{2}(I)\to\mathcal{B}^{2}(I)$ satisfying $\mathcal{N}(C_{B^{2},X}^{1/2})\cap\mathcal{N}(L)=\{0\}$, where $\mathcal{N}(\cdot)$ is the nullspace of the corresponding operator, we define the regularized standardization of $X\in\mathcal{B}^{2}(I)$ with respect to $L$ as the solution to the optimization problem
$X_{\mathrm{st}}^{\alpha,L}=_{B}\operatorname*{arg\,min}_{Y\in\mathcal{B}^{2}(I)}\left\{\|{C}_{B^{2},X}^{1/2}Y\ominus X\|_{\mathcal{B}}^{2}+\alpha\|LY\|_{\mathcal{B}}^{2}\right\},$ (6)
where $C_{B^{2},X}^{1/2}Y=\bigoplus_{i=1}^{\infty}\lambda_{i}^{1/2}\langle\zeta_{i},Y\rangle_{B^{2}}\odot\zeta_{i},$ for $(\lambda_{i},\zeta_{i}),\,i\geq 1$ being the $i$th eigenpair of ${C}_{B^{2},X}$. Observe that $X_{\mathrm{st}}^{\alpha,L}$ is an approximate solution to $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}=X$, where $\alpha$ determines that level of approximation. Locker and Prenter (1980) (Theorem 3.1) guarantees the existence and uniqueness of the solution to (6).
For $L=I$ (the identity operator), the optimization problem (6) can be seen as the Bayes space counterpart of the $\alpha$-standardization in (5). Proposition 3.1 shows the equivalence between (6) in $\mathcal{B}^{2}(I)$ and (5) in $L^{2}(I)$ when applied to the $\mathrm{clr}$-transformed density.
Proposition 3.1
For regularization parameter $\alpha>0$
$X_{\mathrm{st}}^{{\alpha},{I}}=_{B}\operatorname*{arg\,min}_{Y\in\mathcal{B}^{2}(I)}\left\{\|{C}_{B^{2},X}^{1/2}Y\ominus X\|_{\mathcal{B}}^{2}+\alpha\|Y\|_{\mathcal{B}}^{2}\right\}$
if and only if
$\mathrm{clr}(X_{\mathrm{st}}^{{\alpha},I})=\operatorname*{arg\,min}_{Z\in L^{2}(I)}\left\{\|C_{\mathrm{clr},X}^{1/2}Z-\mathrm{clr}(X)\|^{2}+\alpha\|Z\|^{2}\right\},$
where $C_{B^{2},X}$ and $C_{\mathrm{clr},X}$ are the Bayes space covariance of $X$ and covariance of $\mathrm{clr}(X)$, respectively.
Corollary LABEL:cor:L_2_standardization_explicit_form in the Supplement LABEL:supp:aux_results gives the explicit solution to (6) for $L=I$.
Choosing the regularization operator.
Choosing regularization operators to be orthogonal projectors, as discussed in the work of Morigi et al. (2007), can be particularly advantageous when we have prior knowledge of certain features or structures within the data. A key motivation for using orthogonal projection regularization operators is that they allow us to directly control the components of the solution within the range and null space of $L$. By projecting onto a subspace that aligns with known properties of the data, we can tailor the regularization to preserve essential characteristics. The following proposition formalizes these insights and provides further motivation for selecting the appropriate regularization strategy by extending the findings of Morigi et al. (2007) (Theorem 2.3) to optimization (6).
Proposition 3.2
Using the notation and assumptions from Proposition 3.1 let $X_{\mathrm{st}}^{{\alpha},{L}}$ be the unique solution of the minimization problem (6). Consider now the perturbation of the original ill-posed problem $C_{B^{2},X}^{1/2}Y=X$ modifying the right hand side as
$C_{B^{2},X}^{1/2}Y=X\oplus C_{B^{2},X}^{1/2}\tilde{X}$
for some $\tilde{X}\in\mathcal{N}(L)$. Then the unique solution of the associated Tikhonov minimization problem
$\operatorname*{arg\,min}_{Y\in B^{2}(I)}\left\{\|C_{B^{2},X}^{1/2}Y\ominus(X\oplus C_{B^{2},X}^{1/2}\tilde{X})\|_{B^{2}}^{2}+\alpha\|LY\|_{B^{2}}^{2}\right\}$
is given by $X_{\mathrm{1,st}}^{{\alpha},{L}}=X_{\mathrm{st}}^{{\alpha},{L}}\oplus\tilde{X}$.
If we think of $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}$ as a regularized approximation of $X$ (see introductory discussion in Section 3.1 and Berrendero et al. (2020)), we get that $C_{B^{2},X}^{1/2}X_{\mathrm{1,st}}^{{\alpha},{L}}=C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{{\alpha},{L}}\oplus C_{B^{2},X}^{1/2}\tilde{X}$, i.e. the perturbed part was recovered exactly. In other words, Proposition 3.2 suggests that the regularization operator should be chosen so that the known features of the desired solution to (6) can be represented using elements in its null space.
For further illustration, we exemplify the result of Proposition 3.2 by considering a regularization operator based on simplicial functional principal components. For this purpose, observe that $X$ admits a Karhunen-Loève representation
$X=\left(\bigoplus_{i=1}^{k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}\right)\oplus\left(\bigoplus_{i>k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}\right),$
where for $k\geq 0$, such that $\lambda_{k}>\lambda_{k+1}$, $X_{1:k}:=\bigoplus_{i=1}^{k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}$ is the best rank-k approximation of $X$, and $X_{(k+1):\infty}:=\bigoplus_{i>k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}$ the residual. Additionally, it is often argued that the projection of the data onto the first few principal components is an effective way to retain the most significant features of the original dataset (Ramsay and Silverman, 1997; Hron et al., 2016). Consider, therefore, the projection of the data onto the span of the first $k$ eigenfunctions of $C_{B^{2},X}$ to be the data feature we would like to preserve. As $C_{B^{2},X}^{1/2}X_{1:k}\in\mathrm{span}\{\zeta_{1},\dots,\zeta_{k}\}$, by taking $L$ to be the orthogonal projection onto the complement of $\mathrm{span}\{\zeta_{1},\dots,\zeta_{k}\}$, the regularized approximation $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}$ of $X$ is of the form $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}=X_{1:k}\oplus C_{B^{2},X}^{1/2}X_{2,\mathrm{st}}^{\alpha,L}$, where $X_{2,\mathrm{st}}^{\alpha,L}\in\mathrm{span}\{\zeta_{k+1},\zeta_{k+2},\dots\}$ is the $L^{2}$-regularized standardization of $X_{(k+1):\infty}$.
Thus, denoting W(k){W}(k) a projection onto the complement of the space spanned by the first $k\geq 0$ eigenfunctions of $C_{B^{2},X}$, hereinafter we consider standardization of the form
$X_{\mathrm{st},C_{B^{2},X}}^{\alpha,k}=\operatorname*{arg\,min}_{Y\in\mathcal{B}^{2}(I)}\left\{\|{C}_{B^{2},X}^{1/2}Y\ominus X\|_{B^{2}}^{2}+\alpha\|W(k)Y\|_{B^{2}}^{2}\right\},$ (7)
where we add the additional subscript $C_{B^{2},X}$ to emphasize that the regularized standardization is performed with respect to $C_{B^{2},X}$, and superscript $k$ is used instead of $W(k)$ for simplicity. The following proposition gives the exact form of the solution $X_{\mathrm{st},C_{B^{2},X}}^{\alpha,k}$ and also shows that $\mathrm{clr}(X_{\mathrm{st},C_{B^{2},X}}^{\alpha,k})$ is a solution to the $L^{2}$-equivalent of optimization problem (7).
Proposition 3.3
Let $X\in\mathcal{B}^{2}(I)$ be a constant mean random density with Bayes-covariance $C_{B^{2},X}$ and let $(\lambda_{i},\zeta_{i})$ be its $i$th eigenpair; $C_{B^{2},X}\zeta_{i}=\lambda_{i}\odot\zeta_{i}$, $i\geq 1$ . For $k\geq 0$, such that $\lambda_{k}>\lambda_{k+1}$ let $W(k)$ be the projector onto the complement of the space spanned by the first $k$ eigenfunctions $\zeta_{1},\dots,\zeta_{k}$ of $C_{B^{2},X}$. Then there exists a unique solution $X_{\mathrm{st},C_{B^{2},X}}^{{\alpha},k}$ of (7). Additionally,
$X_{\mathrm{st},C_{B^{2},X}}^{{\alpha},k}=_{B}\operatorname*{arg\,min}_{Y\in\mathcal{B}^{2}(I)}\left\{\|{C}_{B^{2},X}^{1/2}Y\ominus X\|_{\mathcal{B}}^{2}+\alpha\|W(k)Y\|_{\mathcal{B}}^{2}\right\}$
if and only if
$\mathrm{clr}(X_{\mathrm{st},C_{B^{2},X}}^{{\alpha},k})=\operatorname*{arg\,min}_{Z\in L^{2}(I)}\left\{\|C_{\mathrm{clr},X}^{1/2}Z-\mathrm{clr}(X)\|^{2}+\alpha\|V(k)Z\|^{2}\right\},$
where $V(k)$ is the projector of the orthogonal complement onto the space spanned by the first $k$ eigenfunctions of the covariance $C_{\mathrm{clr},X}$ of $\mathrm{clr}(X)$. In that case,
$X_{\mathrm{st},C_{B^{2},X}}^{{\alpha},k}=_{B}\sum_{i=1}^{k}{\frac{1}{\lambda_{i}^{1/2}}}\langle X,\zeta_{i}\rangle_{B^{2}}\zeta_{i}\bigoplus\sum_{i=k+1}^{\infty}{\frac{\lambda_{i}^{1/2}}{\lambda_{i}+\alpha}}\langle X,\zeta_{i}\rangle_{B^{2}}\zeta_{i}.$
The technical assumption $\lambda_{k}>\lambda_{k+1}$ ensures well-definiteness of the projector onto the space spanned by the first $k$ eigenfunctions and poses no restriction in practice. Specifically, those PCs having the same variance are considered to be equally important, and should then all or none be regularized.
3.2 Regularized Mahalanobis distance for density data
Having introduced the regularized standardization of data in Bayes space, we are ready to define the corresponding regularized Mahalanobis distance for probability density functions.
Definition 3.1 (Regularized Mahalanobis distance between two densities)
Given a constant $\alpha>0$, and the number of non-regularized components $k\geq 0$, the squared RDMD between $X,Y\in\mathcal{B}^{2}(I)$, with respect to the covariance operator $C$ is given by
[…]
where $X_{\mathrm{st},C}^{{\alpha},k},\,Y_{\mathrm{st},C}^{{\alpha},k}$ are the solutions to
$\min_{Z\in\mathcal{B}^{2}(I)}\left\{\|{C}^{1/2}Z\ominus X\|_{\mathcal{B}}^{2}+\alpha\|W(k)Z\|_{\mathcal{B}}^{2}\right\}\quad\text{and}\quad\min_{Z\in\mathcal{B}^{2}(I)}\left\{\|{C}^{1/2}Z\ominus Y\|_{\mathcal{B}}^{2}+\alpha\|W(k)Z\|_{\mathcal{B}}^{2}\right\},$
respectively.
The RDMD of random density $X$ is, as in the multivariate setting, defined as the RDMD 3.1 between $X$ and its mean $\mu_{B^{2},X}$, with respect to its covariance $C_{B^{2},X}$. For the sake of further referencing, the RDMD of $X\in\mathcal{B}^{2}(I)$ is formally defined as follows.
Definition 3.2 (Regularized Mahalanobis distance for density data)
Given a constant $\alpha>0$, and the number of non-regularized components $k\geq 0$, the squared RDMD of the random density $X\in\mathcal{B}^{2}(I)$ with mean $\mu_{B^{2},X}$ and covariance $C_{B^{2},X}$ is given by
$M^{2}_{\alpha,k,B^{2}}(X)=$ $~M^{2}_{\alpha,k,B^{2}}(X,\mu_{B^{2},X};C_{B^{2},X}),$
where $M^{2}_{\alpha,k,B^{2}}(X,\mu_{B^{2},X};C_{B^{2},X})$ is defined in Definition 3.1. respectively.
One can show that the distribution of the squared RDMD under the assumption of Gaussianity is an infinite weighted sum of independent $\chi^{2}(1)$ variables. For the sake of conciseness, the formal result is given in Proposition LABEL:prop:Mah_distribution in Supplement LABEL:supp:aux_results.
We conclude this section with a discussion of the connection to the functional MD introduced in Galeano et al. (2015) and Berrendero et al. (2020).
Proposition 3.4
Let $X\in\mathcal{B}^{2}(I)$ with w.l.o.g. constant mean $\mu_{B^{2},X}\equiv_{B^{2}}1$ and $\mathrm{clr}(X)\in L^{2}(I)$ be the corresponding clr transformation with covariance $C_{\mathrm{clr}(X)}=\sum_{i=1}^{\infty}\lambda_{i}\xi_{i}\otimes\xi_{i}$, where $(\lambda_{i},\xi_{i})$ is the $i$th eigenpair of $C_{\mathrm{clr}(X)}$, such that $\lambda_{i}>0$ for $i\leq k$. Let further $M^{2}(\mathrm{clr}(X);k)=\sum_{i=1}^{k}\lambda_{i}^{-1}\langle\mathrm{clr}(X),\xi_{i}\rangle^{2}$, and $M_{\alpha}^{2}(\mathrm{clr}(X))=\sum_{i=1}^{\infty}\frac{\lambda_{i}}{(\lambda_{i}+\alpha)^{2}}\langle\mathrm{clr}(X),\xi_{i}\rangle^{2}$ be the k-truncated Mahalanobis distance (Galeano et al., 2015, Definition 2.1) and $\alpha$-Mahalanobis distance (Berrendero et al., 2020, Definition 1), respectively. The following identities then hold:
  • i)
    $M_{\alpha,0,B^{2}}^{2}(X)=M_{\alpha}^{2}(\mathrm{clr}(X))$,
  • ii)
    $\lim_{\alpha\to\infty}M_{\alpha,k,B^{2}}^{2}(X)=M^{2}(\mathrm{clr}(X);k)$.
Proposition 3.4 shows that functional MD introduced by Galeano et al. (2015); Berrendero et al. (2020) can be regarded as extreme cases of the RDMD from Definition 3.2.
4 Robust density PCA (RDPCA)
4.1 Robust Bayes covariance estimation
Given the data sample $X_{1},\dots,X_{n}$ and a subset $H\subseteq\{1,\dots,n\}$, $|H|=h\in[n/2,n]$ we define the trimmed Bayes mean and covariance estimators calculated using an $H$-subset of the data as
$\hat{\mu}_{B^{2},H}=\frac{1}{h}\odot\bigoplus_{i\in H}X_{i},\quad\hat{C}_{B^{2},H,c_{H}}=c_{H}\frac{1}{h}\odot\bigoplus_{i\in H}(X_{i}\ominus\hat{\mu}_{B^{2},H})\otimes_{B^{2}}(X_{i}\ominus\hat{\mu}_{B^{2},H}).$ (8)
The constant $c_{H}$ is a scaling factor to adjust for trimming and is determined under the assumption of Gaussianity. For that purpose, we use the corollary of Proposition LABEL:prop:Mah_distribution.
Corollary 4.1
Let $X\in\mathcal{B}^{2}(I)$ be a random density such that $\mathrm{clr}(X)\in L^{2}(I)$ is a Gaussian random process, and let $\hat{C}_{B^{2},X}$ and $\hat{\mu}_{B^{2},X}$ be strongly consistent estimators of the covariance operator $C_{B^{2},X}$ and the constant mean $\mu_{B^{2},X}\equiv_{B^{2}}1$, respectively. Then, for $\alpha>0$ and $k\in\mathbb{N}_{0}$, such that $\lambda_{k}>\lambda_{k+1}$, $M_{\alpha,k,B^{2}}^{2}(X,\hat{\mu}_{B^{2},X};\hat{C}_{B^{2},X})$ converges in distribution to
$M^{2}_{\alpha,k,B^{2}}(X)\sim\begin{cases}y+\sum_{i=k+1}^{\infty}\frac{\lambda_{i}^{2}}{(\lambda_{i}+\alpha)^{2}}\eta_{i},\text{ for }k\geq 1,\\ \sum_{i=1}^{\infty}\frac{\lambda_{i}^{2}}{(\lambda_{i}+\alpha)^{2}}\eta_{i},\text{ for }k=0,\end{cases}$
where $\eta_{i}\sim\chi^{2}(1)$, $i\geq 1$ are mutually independent and also independent of $y\sim\chi^{2}(k)$, and $\lambda_{i}$, $i\geq 1$ is the $i$th eigenpair of $C_{B^{2},X}$.
For more details on the necessity of scaling the trimmed covariance, see, e.g., Boudt et al. (2020); Oguamalam et al. (2024). Using the notion of RDMD as a centrality measure, we choose the trimming subset $H_{0}$ as a solution to
$H_{0}$ $=\operatorname*{arg\,min}_{\{H\subset\{1,\dots,n\};|H|=h\}}\frac{1}{h}\sum_{i\in H}M_{\alpha,k,B^{2}}^{2}(X_{i},\mu_{B^{2},X};C_{B^{2},X}),$ (9)
where $\alpha>0$ and $k\in\mathbb{N}_{0}$ are a priori chosen regularization parameters.
Since the true mean and covariance operator are unknown, we replace them in (9) with the current robust estimates. This approach leads to an implicit equation for obtaining the optimal subset:
$H_{0}$ $=\operatorname*{arg\,min}_{\{H\subset\{1,\dots,n\};|H|=h\}}\frac{1}{h}\sum_{i\in H}M_{\alpha,k,B^{2}}^{2}(X_{i},\hat{\mu}_{B^{2},H_{0}};\hat{C}_{B^{2},H_{0},c_{H_{0}}}).$ (10)
(Oguamalam et al., 2024, Lemma 3.1) guarantees that the equation (10) has a solution.
Finally, Corollary 4.1 implies that $c_{H_{0}}$ can be estimated by matching the sample median of $\{M_{\alpha,k,B^{2}}^{2}(X_{i},\hat{\mu}_{B^{2},H_{0}};\hat{C}_{B^{2},H_{0},c_{H_{0}}}),\,i=1,\dots,n\}$ with the estimate of the median of the corresponding limiting distribution given in Corollary 4.1, where the eigenvalues $\lambda_{i},\,i\geq 1$ of $C_{B^{2},X}$ are again replaced by their robust estimates;
$c_{H_{0}}=\mathrm{med}\left\{M_{\alpha,k,B^{2}}^{2}(X_{i},\hat{\mu}_{B^{2},H_{0}};\hat{C}_{B^{2},H_{0},c_{H_{0}}}),\,i=1,\ldots,n\right\}/\mathrm{med}\left\{y+\sum_{i=k+1}^{\infty}\frac{\hat{\lambda}_{i}^{2}}{(\hat{\lambda}_{i}+\alpha)^{2}}\eta_{i}\right\},$
where $\eta_{i}\sim\chi^{2}(1)$, $i\geq 1$, are mutually independent and are independent of $y\sim\chi^{2}(k)$ for $k\geq 1$. We note that, for the purpose of PCA, the scaling is redundant. We include it, however, for completeness.
Remark 4.1
Corollary 4.1 can be used to detect potential outliers by comparing the calculated RDMD to an appropriate quantile from the limiting reference distribution; see Remark 2.1. However, we emphasize how RDMD and the underlying robust estimation do not rely on Gaussianity. In practice, RDMD values can highlight atypical observations, with non-parametric cutoffs, like, e.g., empirical quantiles or boxplots, used without distributional assumptions. The Gaussianity is invoked only to enable an automated, theoretically grounded cutoff for outlier detection, which may be useful in applications such as re-weighting in robust covariance estimation, as e.g. in Rousseeuw and Driessen (1999); Berrendero et al. (2020); Oguamalam et al. (2024).
While general consistency results under a non-contaminated model like those of Croux and Haesbroeck (1999); Butler et al. (1993) for the MCD are beyond the scope of this work, we provide a partial result under a clean model, where the trimming proportion grows sublinearly with the sample size. In this setting, we show that the robust estimators (8) and the associated principal functions converge to the population counterparts. These results are given in Lemma LABEL:sup_lemma:consistency in Supplement Section LABEL:supp:aux_results, along with simulation studies that empirically support this convergence in broader scenarios; see Figures LABEL:fig:additionalResults_consistency_sim1 and LABEL:fig:additionalResults_consistency_sim2.
4.2 Algorithm
We solve (10) iteratively, replacing the RDMD in the $j$th iteration with its estimate using the (j1)(j-1)st optimal subset:
$H_{j}=\operatorname*{arg\,min}_{\{H\subset\{1,\dots,n\};|H|=h\}}\frac{1}{h}\sum_{i\in H}\hat{M}_{\alpha,k,B^{2}}^{2}(X_{i},\hat{\mu}_{B^{2},H_{j-1}};\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}),$ (11)
where $\hat{\mu}_{B^{2},H_{j-1}}$ and $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$ are the trimmed sample mean and covariance operator based on the subset $H_{j-1}$, respectively. Thus, in the $j$th step of the procedure, we identify $h$ observations with the smallest regularized Mahalanobis distances to the robust mean estimate $\hat{\mu}_{B^{2},H_{j-1}}$, w.r.t. the current robust covariance estimate $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$. However, in practice, we don’t directly observe functions, but rather discrete realizations of the underlying functional observations, or datasets from which such functions are estimated, which is often the case with probability density functions. Let therefore
$X_{i}(t):t\in\{t_{1},\dots,t_{p}\},\,i=1,\dots,n,$ (12)
represent such a discretely observed sample. When the grid is sufficiently dense, the functional inner products required in computing the RDMD can be accurately approximated by integral sums; see, e.g., Ostebee and Zorn (2002). Accordingly, both $X_{i}$ and related quantities such as $\hat{\mu}_{B^{2},H}$ are treated as vectors, and the sample covariance operator $\hat{C}_{B^{2},H,c_{H}}$ is replaced by a matrix of pointwise covariances. The pseudocode for obtaining the robust pointwise estimators of the mean, covariance, and principal functions is given in Algorithm LABEL:alg:alg1 in Supplement Section LABEL:supp:aux_results. After obtaining robust estimates on the observed grid, one may interpolate these estimates to obtain a full representation over the domain, though this step lies outside the scope of this work.
Alternatively, one can first convert discretized observations into a functional representation by projecting them onto a suitable basis. This enables evaluating each function at any point $t\in I$, not just at the sampled time points. Using a fixed number $M\in\mathbb{N}$ of chosen basis functions $\phi_{1}(t),\ldots,\phi_{M}(t):I\rightarrow\mathbb{R}^{+},j=1,\ldots,M,$ each observation can be approximated as:
$X_{i}^{M}(t)=\bigoplus_{j=1}^{M}C_{i,j}\odot\phi_{j}(t),\quad i=1,\dots,n,$ (13)
where $C_{i,j}]\in\mathbb{R}^{n\times M}$ is the matrix of basis coefficients, and the number of basis functions $M$ is typically chosen to strike a balance between accurately capturing the observed data and enforcing smoothness to mitigate noise. While this representation is exact for functions of finite rank, it can also serve as an additional smoothing technique when the basis and the number of components are appropriately chosen. An (orthonormal) basis of our choice is the compositional spline basis introduced in Machalová et al. (2021), designed specifically for the representation of the relative functional data, thus satisfying the unit integral constraint.
Proposition 4.1 shows that the RDMD can be computed exactly in this framework, using only the matrix of basis coefficients $C_{i,j}]$. The corresponding routine for robust covariance and SPCA estimation is given in Algorithm 1.
Proposition 4.1
Let $\boldsymbol{\Phi}(t)=(\phi_{1}(t),\dots,\phi_{M}(t))^{\prime}$, $t\in I$ be an orthonormal basis, and let $X_{i},i=1,\dots,n$ admit a rank-M representation as in (13), where $\textbf{C}=[C_{i,j}]\in\mathbb{R}^{n\times M}$ is the the matrix of basis coefficients. Then, for $H\subseteq\{1,\dots,n\}$, $|H|=h,$ and parameters $\alpha>0$, $k\in\mathbb{N}_{0}$ $\hat{M}^{2}_{\alpha,H,B^{2}}(X_{i},\hat{\mu}_{B^{2},H};\hat{C}_{B^{2},H,c_{H}})=(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H})^{\prime}\textbf{C}\textbf{C}_{H}(\textbf{C}_{H}+\alpha/c_{H}\textbf{W}(k))^{-2}\textbf{C}\,^{\prime}(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H}),$ where $\textbf{C}_{H}=\frac{1}{h}\textbf{C}\,^{\prime}(\textbf{I}_{n,H}-\frac{1}{h}\textbf{J}_{n,H})\textbf{C}\in\mathbb{R}^{M\times M}$, $\mathbf{1}_{H}\in\mathbb{R}^{n}$ is such that $(\mathbf{1}_{H})_{i}=1$ if $i\in H$ and 0 otherwise, $\textbf{J}_{n,H}=\mathbf{1}_{H}\mathbf{1}_{H}^{\prime}$, and $\textbf{I}_{n,H}=\mathrm{diag}(\mathbf{1}_{n,H})\in\mathbb{R}^{n\times n}$ is a diagonal matrix with $\mathbf{1}_{H}$ on its diagonal. $\textbf{W}(k)$ is a projection matrix onto the space spanned by the span of the first $k$ eigenvectors of $\textbf{C}_{H}$; if $k=0$, $\textbf{W}(k)=\textbf{I}_{M}.$
Remark 4.2
Our framework assumes some level of smoothness in the underlying functional data, an assumption embedded via regularization. In practice, this means that the methodology can also be justifiably applied to data observed on a dense, regular grid without requiring basis expansion. In fact, pre-smoothing or excessive preprocessing may mask outlying behavior, potentially interfering with outlier detection (Hubert et al., 2015). Therefore, both discretized and functional representations are supported within our framework—the choice between them should depend on the nature of the data analyst’s preference.
Input : Orthonormal basis $\boldsymbol{\Phi}=(\phi_{1},\dots,\phi_{M})\in\mathcal{B}^{2}$, sample ${X}_{1},\dots,{X}_{n}\in\mathcal{B}^{2}$, subset size $h$, initial subset $H_{1}\subset\{1,\dots,n\}$ with $|H_{1}|=h$, regularization parameter $\alpha>0$, projection dimension $k\leq n$, and tolerance level $\varepsilon_{k}>0$;
1
[…]
3 Express $Y_{i}$ in basis $\tilde{\boldsymbol{\Phi}}$ as $Y_{i}=\sum_{j=1}^{M}C_{i,j}\tilde{\phi}_{j}$, and collect $\boldsymbol{C}=[C_{i,j}]\in\mathbb{R}^{n\times M}$;
4 do
5 $H_{0}\leftarrow H_{1}$;
6 $\boldsymbol{C}_{H_{0}}=\frac{1}{h}\boldsymbol{C}^{\prime}(\textbf{I}_{n,H_{0}}-\frac{1}{h}\textbf{J}_{n,H_{0}})\boldsymbol{C}=\sum_{i=1}^{M}{\hat{\lambda}}_{i,H_{0}}\textbf{u}_{i}\textbf{u}_{i}^{\prime}$;
7   Set initial scaling parameter $c_{1,H_{0}}\leftarrow 1$;
8 do
9    $c_{0,H_{0}}\leftarrow c_{1,H_{0}}$;
10      Set $\tilde{V}(k)\leftarrow I-\sum_{j=1}^{k}\hat{\lambda}_{j,H_{0}}\boldsymbol{u}_{j}\boldsymbol{u}_{j}^{\prime}$;
11    $d_{i,H_{0},c_{0}}^{2}\leftarrow(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H_{0}})^{\prime}\boldsymbol{C}\boldsymbol{C}_{H_{0}}(\boldsymbol{C}_{H_{0}}+\alpha/c_{0,H_{0}}\tilde{V}(k))^{-2}\boldsymbol{C}^{\prime}(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H_{0}})$;
12      Generate $\eta_{k+1},\ldots,\eta_{n}\sim\chi^{2}(1)$, $y\sim\chi^{2}(k)$;
13    $c_{1,H_{0}}\leftarrow{\mathrm{med}\left\{d_{i,H_{0},c_{0}}^{2}:i=1,\dots,n\right\}}/{\mathrm{med}\left\{y+\sum_{i=k+1}^{M}\frac{\hat{\lambda}_{i,H_{0}}^{2}}{(\hat{\lambda}_{i,H_{0}}+\alpha/c_{0,H_{0}})^{2}}\eta_{i}\right\}}$;
14 while $(c_{1,H_{0}}-c_{0,H_{0}})^{2}\geq\varepsilon_{k}$;
15  Order $d^{2}_{(i_{1},H_{0},c_{1})}\leq\dots\leq d^{2}_{(i_{n},H_{0},c_{1})}$;
16   Set $H_{1}\leftarrow\{i_{1},\dots,i_{h}\}$;
17
18while $H_{0}\neq H_{1}$;
19Solve symmetric eigenproblem $c_{1,H_{1}}\boldsymbol{C}_{H_{1}}(u_{j})=\hat{\lambda}_{j}u_{j};$
20 Recalculate $\xi_{j}=u_{j}^{\prime}\tilde{\boldsymbol{\Phi}}$ and set $\hat{\xi}_{j,H_{\mathrm{opt}}}\leftarrow\xi_{j},\hat{\zeta}_{j,H_{\mathrm{opt}}}\leftarrow\mathrm{clr}^{-1}(\xi_{j})$ and $\hat{\lambda}_{j,H_{\mathrm{opt}}}\leftarrow\hat{\lambda}$;
Output : Optimal subset $H_{\mathrm{opt}}:=H_{1}$, scaling factor $c_{H_{0}}$, squared robust RDMD $(c_{H_{0}}^{-1}d_{1,H_{0},c_{1}}^{2},\dots c_{H_{0}}^{-1}d_{n,H_{0},c_{1}}^{2})$, robust eigenvalues $\{\hat{\lambda}_{j,H_{\mathrm{opt}}}\}_{j\geq 1}$ and corresponding robust FPCs $\{\hat{\zeta}_{j,H_{\mathrm{opt}}}\}_{j\geq 1},\{\hat{\xi}_{j,H_{\mathrm{opt}}}\}_{j\geq 1}$ in Bayes and clr space, respectively.
Algorithm 1 RDPCA for data expressed in a fixed basis
Selecting the regularization parameters
The regularized Mahalanobis distance depends on the tuning parameters $\alpha>0$ and $k\in\mathbb{N}_{0}$. The choice of $\alpha$ balances the smoothness of the regularized data with how well it fits the observed data. If $\alpha$ is too large, over-regularization will result in poor fit, with large residuals $\|\hat{C}_{B^{2},H_{\mathrm{opt}}}^{1/2}(X_{i,\mathrm{st},\hat{C}_{B^{2},H_{\mathrm{opt}}}}^{{\alpha},k})\ominus X_{i})\|_{B^{2}}^{2}$, $i=1,\ldots,n$. Conversely, if $\alpha$ is too small, the solution will overfit data errors, leading to large values of $\|\hat{W}(k)X_{i,\mathrm{st},\hat{C}_{B^{2},H_{\mathrm{opt}}}}^{{\alpha},k}\|^{2}$. To determine $\alpha$, we adapt the approach from Oguamalam et al. (2024), iteratively adjusting $\alpha$ so that the Bayes covariance of the standardized data has eigenvalues close to zero (for the latter part of the spectrum) or some fixed constant (for the signal part), thus maximizing the signal-to-noise ratio.
On the other hand, we can think of $k$ parameter that prevents important signal features from being over-smoothed. Hron et al. (2016) argues how for densities from the $k_{0}$ - exponential family (with $k_{0}$ parameters), the number of principal components belonging to strictly positive eigenvalues is precisely $k_{0}$. E.g., for the Gaussian family with the known mean $k_{0}=1$; see Hron et al. (2016) for more details. Therefore, in Section 5 we fix $k=1$. In more complex situations where the underlying distribution is unknown, selecting $k$ such that the first $k$ principal functions contain at least $90\%$ of variability seems reasonable. This rule of thumb will be applied in Section 6.
Finally, in the supplement, we provide some ideas about the choice of the subset size $h$.
5 Simulations
Within two simulation studies, we evaluate the accuracy of the proposed robust approach in estimation of the covariance and principal components under contamination with a fixed proportion $c\in\{0,0.05,0.1,0.2\}$ of the data is replaced by outliers. Outliers are introduced via Huber’s $\varepsilon$-contamination model Huber (1964) as $X\sim(1-c)F_{reg}+cF_{out}$, where $F_{reg}$, $F_{out}$ denote the distribution of the regular and outlying observations, respectively.
To mitigate the effect of the data processing on the results, in the simulations we use the raw data sampled on a dense grid and robustly estimate the covariance and PCs using Algorithm LABEL:alg:alg1 from the supplement, with the regularization parameter $\alpha$ automatically chosen according to the procedure discussed in Section 4.2. Furthermore, we fix $h=\lfloor 0.75n\rfloor$ and $k=1$.
We assess both the accuracy and robustness of the covariance operator and principal component estimation, and compare the results with those of the non-robust SFPCA method. To measure the performance of RDPCA, we consider two statistics: the integrated square error (ISE), approximating the norm of the difference between estimates of the true and sub-sampled covariances, and the mean cosine between the first five pairs of eigenfunctions based on the same estimates. For more details, see Supplement Section LABEL:supp:simulations. For completeness, we also examine the performance of several outlier detection methods applied after the clr transformation, as commonly done in the functional data analysis literature (Arribas-Gil and Romo, 2014; Berrendero et al., 2020; Oguamalam et al., 2024). Detailed results are provided in Supplementary Section LABEL:sec:add_simul_results_supp.
5.1 Simulation 1
In the first simulation, in each of $100$ replications, we consider a data set of size $n=500$ consisting of $n-\lfloor cn\rfloor$ regular- and $\lfloor cn\rfloor$ contaminated curves, $c\in\{0,\,0.05,\,0.1,\,0.2\}$. Each curve is estimated using kernel density estimation (KDE) on a generated sample and evaluated at an equidistant grid of $p=50$ time points. The construction of a regular/outlying sample is as follows:
$X_{\mathrm{reg}}=~\mathrm{KDE}\left(Y_{1},\ldots,Y_{N}\right),~X_{\mathrm{out}}=~\mathrm{KDE}\left(Y_{1},\ldots,Y_{N},\tilde{Y}_{1},\ldots,\tilde{Y}_{\lfloor 0.1N\rfloor}\right)$
where $Y_{i}\sim N(0,1),~\tilde{Y}_{i}\sim U([q_{1};q_{2}]\cup[-q_{2};-q_{1}]),N=250$ and $q_{1},q_{2}$ being the $0.1\%$ and $0.5\%$ quantile of the standard normal distribution, respectively. KDE was calculated using the R-package kde1d (Nagler, 2024).
By the construction of these densities, a fixed proportion $c$ of the data is contaminated within the tails of the distribution. Figure 1 presents an example of this model, displaying both the densities and their transformed counterparts, for a contamination rate of $c=0.2$. In the supplement, we extend this simulation with two additional models, where the densities are constructed in a similar fashion for different distributions of $Y_{i}$ and $\tilde{Y}_{i}$.
Figure 2 depicts the accuracy of the estimation of principal components w.r.t. the contamination rate $c$. Both the ISE and the mean cosine between the first five pairs of principal functions suggest that RDPCA significantly outperforms SFPCA in the presence of outliers. Interestingly, the right panel of Figure 2 indicates that RDPCA outperforms SFPCA in estimating the principal functions even in the non-contaminated setting.
Figure 1: Visualization of density functions (left) and clr transformed counterparts (right). Solid curves represent the main processes, while the dashed ones indicate the outliers.
Figure 2: Mean ($\pm$ standard error) ISE between the estimated and true covariances (left) and cosine similarity between the first five pairs of estimated and true eigenfunctions (right), obtained using RDPCA (solid) and SFPCA (dashed).
Figure 3 illustrates how the robust correlation function nicely aligns with the correlation function based on a Monte Carlo estimate of the true underlying model. On the contrary, the non-robust correlation reveals a significantly different structure. In Figure 4 we compare the squared RDMD computed using RDPCA to the squared RDMD obtained using SPCA and the squared robust $\alpha$-Mahalanobis (Berrendero et al., 2020) distance, ignoring the compositional structure of the densities. The dashed lines indicate the $95\%$ quantile of the limiting distribution of these distances assuming Gaussianity and estimated by Monte Carlo simulation; see Corollary 4.1. A density is considered outlying if its squared Mahalanobis distance exceeds this cutoff. The circles represent the true regular observations, whereas the triangles correspond to the outliers. According to the cutoff values (dashed lines), only RDPCA can determine outliers in a meaningful way. In the supplement, we added some additional results related to this comparison.
Figure 3: True (left), RDPCA with $\alpha=0.32$ (middle), and SFPCA (right) correlation function, for the example shown in Figure 1.
Figure 4: Distance-distance plot of squared robust vs. non-robust RDMD (left) and robust $\alpha$-Mahalanobis (Berrendero et al., 2020) (right) distances for example of Figure 1. The regularization parameter is $\alpha=0.32$. Dashed lines indicate the corresponding cutoff values under Gaussianity. Circles and triangles correspond to true regular and outlying observations, respectively.
5.2 Simulation 2
In the second simulation, we generate data sets of size $n=500$ according to
$\mathrm{clr}(X_{\mathrm{reg}}(t))=$ $~s_{1}\xi_{1}+s_{2}\xi_{2}+s_{3}\xi_{3}+s_{4}\xi_{4},$ (14)
$\mathrm{clr}(X_{\mathrm{out}}(t))=$ $~s_{1}\xi_{1}+s_{2}\xi_{2}+s_{3}\xi_{3}+s_{4}\xi_{4}+s_{5}\xi_{5},$ (15)
where $\xi_{1}=\sqrt{2}\mathrm{sin}(2\pi t),\xi_{2}=\sqrt{2}\mathrm{cos}(2\pi t),\xi_{3}=\sqrt{2}\mathrm{sin}(4\pi t),\xi_{4}=\sqrt{2}\mathrm{cos}(4\pi t)$ are the eigenfunctions of the covariance operator of the clr transformed densities. The scores $s_{1},s_{2},s_{3},s_{4}$ are drawn, in one setting, from a normal distribution, $s_{i}\sim N(0,\lambda_{i})$ or, in another setting, from a multivariate t-distribution; $(s_{1},s_{2},s_{3},s_{4})\sim t_{5}(\boldsymbol{0},\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}))$ and $(s_{1},s_{2},s_{3},s_{4},s_{5})\sim t_{5}(\boldsymbol{0},\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}))$. In both cases we set $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5})=(2,1,1/2,1/4,4)$. The resulting process is then evaluated at an equidistant grid of $p=100$ time points. Visualization of the data is given in the supplement in Figure LABEL:fig:dataClrSim2Mod12.
Each setting is replicated 100 times, and the accuracy of RDPCA is evaluated by the ISE of the difference between the estimated and the true covariance, as well as the mean cosine between the first four pairs of eigenfunctions based on the estimated and the true covariance. The results are shown in Figure 5, and again demonstrate that the robust procedure outperforms SFPCA in the presence of outliers. Interestingly, for $t$-distributed scores, even in the setting with no outliers, the robust procedure yields better results.
Figure 5: Mean ($\pm$ standard error) ISE between the estimated and true covariances (left) and cosine similarity between the first five pairs of estimated and true eigenfunctions (right), obtained using RDPCA (solid) and SFPCA (dashed). The notation “normal" and “t" refer to the distribution of the scores in the underlying model as given at the beginning of the section.
Finally, we compare the estimated ratios of explained variance from the robust and non-robust methods to the true values. As shown in Figure 6, the robust method provides estimates that closely match the true ratios.
Figure 6: Estimated ratio of explained variance for the first five principal functions over different score distributions at $c=0.2$. The dashed reference lines depict the true proportion. The notation “normal" and “t" refer to the distribution of the scores in the underlying model as described at the beginning of this section.
6 Real-data example
To illustrate the practical performance of RDPCA, we apply it to two real-data examples. The computations were carried out on a machine equipped with an AMD Ryzen 9 5900HX core with Radeon Graphics (16 cores, 3.30 GHz) and 32 GB of RAM, running Windows 11 Pro. The implementation was written in R (version 4.4.2). Calculations connected to RDPCA were run using 15 parallel cores and required approximately 5 and 1 minute(s) for the first and second example, respectively.
In addition to robust PCA, during the examples, we also examine atypical curves according to Algorithm 1. For this, we estimate the $95\%$-quantile of the limiting distribution of the RDMD under the assumption of Gaussianity; see Corollary 4.1. A density is then considered outlying if its squared RDMD exceeds this cutoff.
6.1 EPXMA spectra
The considered dataset consists of electron probe X-ray microanalysis (EPXMA) spectra of various glass vessel types. This type of analysis is commonly used to determine the chemical composition of glasses by measuring X-rays at different wavelengths emitted from a sample after being hit by a beam of electrons. As is commonly done in the analysis of spectral data, we analyze EPXMA spectra as relative functional data.
The dataset contains 180 such observations divided into four distinct groups corresponding to different glass vessel types. The largest group, referred to as “sodic", comprises 145 spectra. The remaining groups, designated as “potassic", “potasso-calcic", and “calcic", are represented by $15,10,$ and $10$ spectra, respectively. It is important to note that the latter 38 spectra within the sodic group were acquired under varying experimental conditions. Each observation is characterized by measurements across 1920 wavelengths. For the purposes of this analysis, we focus on wavelengths 51–750 as these encompass the primary variability observed in the data and to avoid measurement artifacts; for more details, see Lemberge et al. (2000). This results in n=180,p=700n=180,\,p=700.
We analyze the data using Algorithm 1, where the data is expressed using $31$ compositional splines. The left plot of Figure 7 shows the smoothed trajectories of the clr transformed observations colored according to their respective glass vessel type. For the analysis, we consider the first 107 observations from the sodic group as regular observations. We use $h=\lceil 0.5n\rceil$ and $k=4$, and automatically select $\alpha=0.03$; see Section 4.2.
The right plot of Figure 7 displays the squared robust versus the non-robust RDMD. The non-robust approach detects virtually no outlying observations, while the robust procedure can clearly differentiate between the majority of “sodic" and the remaining observations.
Figure 7: Left: clr transformed spectra colored according to glass vessel type. Right: Distance-distance plot of the squared robust vs. non-robust RDMD. Dashed lines indicate the theoretical cutoff values under Gaussianity.
Figure 8 showcases the consequent influence onto the covariance, where we compare smoothed robust to non-robust correlation function. The non-robust correlation shows its bias towards the outlying curves at parts of the domain where significant differences arise.
Figure 8: Correlation function of EPXMA data based on RDPCA and SFPCA.
6.2 Fertility data
The second data set consists of age-specific fertility rates (ASFR) across different countries. ASFR represents the number of women giving birth per $1,000$ within specific age groups, providing a detailed view of how fertility changes with age. The resulting curves are analyzed as density data because we are interested in the underlying distribution across the different age groups rather than the magnitude of it. We examine data from the Human Fertility Database, available at https://www.humanfertility.org/Home/Index, focusing on the years $1956$, $1990$, and $2019$, to gain insight into the evolution of fertility dynamics.
In this example, we apply Algorithm 1 with 21 compositional splines as basis functions. The data is visualized in Figure 9. Outlying countries, according to their squared RDMD, are represented by dashed curves. During the early stages of this period, peak fertility was primarily achieved by young women in their early 20s. Large families were common, reflected in a rather slow decrease in the majority of the fertility curves. During the following decades, mostly influenced by the introduction of reliable contraceptives and a change in the mindset about life quality, later first pregnancies became more and more common. Throughout Europe, this phenomenon started earlier in the West than in the East, as can be seen by the fertility curves corresponding to the year $1990.$ Since then, the peak fertility has steadily increased and sits now at around 30 years. For more details, we refer to the insightful article by Beaujouan (2020) and references therein.
While the variability in the Bayes space seems to be present within the peaks, the upper row of Figure 9 shows that the main variability within the clr space is significantly higher in the tails of the distribution; see also Hron et al. (2016).
Figure 9: Fertility densities (lower row) and clr transformed densities (upper row) corresponding to years 1956, 1990, and 2019. Curves identified as outlying by Algorithm 1 are displayed as dashed lines.
In Figure 10, the first two principal functions are displayed. In the brackets, we included the corresponding percentage of explained variance. For each year, the first component mainly captures the variability in the tails of the densities. Here, the principal functions focus on the left tail and the following ages until the early 20s. The second principal function contains significantly less information, mainly explaining the right tail and the peak fertility.
Figure 10: Comparison between the first two robust principal functions of the fertility data for years $1956,1990,2019$. Corresponding percentage of explained variance in brackets (left to right).
Following the construction of robust principal components, we look at the corresponding estimated principal component scores $\hat{s}_{j}$ associated with each country. In Figure 11, the first two robust scores for each country are visualized. For 1956, the outliers, as indicated by the triangles, all lay at the border of the data cloud. For Japan, the first and second scores are exceedingly high, indicating not only relatively low fertility in younger age groups but also proportionally very high peak fertility. For the outliers of 1990 and 2019, we obtain a similar picture dominated by mostly extreme scores. For 1990, this mainly reflects countries that have an unusually high or low fertility for younger and older age groups. Similarly, in 2019, these extreme values explain high fertility rates at younger ages as well as late peak fertility.
Figure 11: Robust PC scores for the first two eigenfunctions for the years $1956,1990,2019$. Circles correspond to regular observations, while the triangles indicate outlying countries.
7 Conclusions and discussion
Despite the attention that the Bayes space methodology for density data analysis has attracted over the last decades, little focus has been paid to robust frameworks necessary for meaningful analysis in the presence of anomalies. Therefore, this paper introduces robust density PCA (RDPCA) as a methodology to robustly estimate the covariance operator and functional principal components of densely sampled univariate density data. Within this newly proposed method, we extend the notion of a functional Mahalanobis distance (Berrendero et al., 2020; Galeano et al., 2015) to the Bayes space. The resulting regularized density Mahalanobis distance (RDMD) leads to a robust covariance estimation based on a subsample of those curves that best fit the underlying distribution. Based on this estimate, robust PCs for density function can be determined as summarized in Algorithm LABEL:alg:alg1. If the data can be expressed in a fixed basis, we provide an equivalent approach in Algorithm 1. During simulations, we demonstrated the efficiency of RDPCA as well as the advantages it has against a non-robust approach in the presence of outliers. Furthermore, if outliers arise within the tails of the densities, the clr transformation can amplify these deviations. Especially in these scenarios, we observed that common robust procedures that do not respect the nature of the data would fail.
One of our initial assumptions was that the data had been observed at a dense grid. In the case where the data is only available at a sparse grid, further extensions would require smoothing by an appropriate basis, e.g., CB-splines (Machalová et al., 2021) specifically developed for Bayes spaces. Naturally, the next step from univariate data would be to consider multivariate densities. During multivariate functional data analysis, each observation contains the recording of several “functional" variables. In this setting, not only is the covariance between different time points considered, but also the relation between individual variables. If the second dimension is also continuously observed, the observations are so-called random surfaces. As these statistical fields show increasingly practical relevance (Berrendero et al. (2011); Górecki et al. (2018); Dai and Genton (2018); Masak and Panaretos (2023) and references therein), expanding RDPCA to these areas seems worthwhile. Calculating the RDMD involved regularizing by a suitable operator that smooths out unwanted noise components while keeping the relevant signal within the (uncontaminated) data unaffected. Another class of meaningful operators that utilize the functional nature of the data are differential operators, often used in Tikhonov regularization. One has to keep in mind that these operations have yet to be defined for Bayes spaces. Next to PCA, the regularized Mahalanobis distance could be used for concepts like linear or quadratic discriminant analysis for the classification of density data. As these methods rely on similarity measures involving several covariance operators, the robust classification of densities based on the RDMD would be suitable.
Supplementary material
Supplement: The supplement contains auxiliary results, further developing ideas on RDPCA. Next, it provides additional simulation analysis corresponding to Section 5.1 and 5.2, and a further analysis of the fertility data from Section 6.2. The supplement finishes with the proofs of theoretical results.
Acknowledgement
The authors acknowledge support from the Austrian Science Fund (FWF), project number I 5799-N, and the Czech Science Foundation, project numbers 22-15684L and 25-15447S. The present research has also been partially supported by MUR, grant Dipartimento di Eccellenza 2023-2027.
Disclosure statement
The authors report there are no competing interests to declare.